# Asymptotics of dynamic ASEP using duality

### v1

@@ 2.3.3 Tracy–Widom contour integral formulas
$\begin{array}[]{c}\sum_{\sigma\in\mathbf{S}_{N}}\operatorname{sgn}\sigma\left(\prod_{i<j}\left(\alpha+\beta\xi_{\sigma(i)}\xi_{\sigma(j)}-\xi_{\sigma(i)}\right)\frac{\xi_{\sigma(2)}\xi_{\sigma(3)}^{2}\xi_{\sigma(4)}^{3}\cdots\xi_{\sigma(N)}^{N-1}}{\left(1-\xi_{\sigma(1)}\xi_{\sigma(2)}\xi_{\sigma(3)}\cdots\xi_{\sigma(N)}\right)\cdots\left(1-\xi_{\sigma(N-1)}\xi_{\sigma(N)}\right)\left(1-\xi_{\sigma(N)}\right)}\right)\\ =\alpha^{N(N-1)/2}\frac{\prod_{i<j}\left(\xi_{j}-\xi_{i}\right)}{\prod_{j}\left(1-\xi_{j}\right)}\end{array}$ (2)
@@ 2.5 From duality to asymptotics
$\lim_{L\rightarrow\infty}\sum_{S}\left(\sum_{x\in\mathcal{S}}\left|(n_{\lambda}^{S})^{-1/2}D_{\lambda}^{S}[x]-(n_{\overline{\lambda}}^{S})^{-1/2}D_{\overline{\lambda}}^{S}[x]\right|\cdot\left|Q_{\overline{\lambda}}(L,x)\right|\right)\left(\sum_{y\in\mathcal{S}}\left|h(L)[y](n_{\lambda}^{S})^{-1/2}D_{\lambda}^{S}(y)W_{\lambda}(y)\right|\right)=0$ (8)
@@ 3.3 Particle locations in (multi–species) ASEP
×$\hskip 72.26999pt\times\int_{C_{r}}\cdots\int_{C_{r}}A_{\sigma}\prod_{i=1}^{K}\xi_{\sigma(i)}^{x_{i}-y_{\sigma(i)}-1}\prod_{i=K+1}^{N}\frac{\xi_{\sigma(i)}^{M_{i}-y_{\sigma(i)}-1}}{1-\xi_{\sigma(i)}}e^{\sum_{i=1}^{N}(p\xi_{i}^{-1}+q\xi_{i}-1)t}d\xi_{1}\cdots d\xi_{N}.$ (11)
@@ 4.2.2 Proof of Proposition 2
$\sum_{\sigma\in S^{+}_{k}(N)}\left[\prod_{j=1}^{k-1}\xi_{\sigma(j)}^{\bar{x}_{j}-y_{\sigma(j)}-1}\right]\frac{\xi_{\sigma(k)}^{-M-y_{\sigma(k)}-1}\xi_{\sigma(k+1)}^{-M-y_{\sigma(k+1)}-1}}{(1-\xi_{\sigma(k)})(1-\xi_{\sigma(k+1)})}\left[\prod_{j=k+2}^{N}\xi_{\sigma(j)}^{\bar{x}_{j}-y_{\sigma(j)}-1}\right]\\ \times\alpha^{-1}\left(A_{\sigma}(\alpha+\beta\xi_{\sigma(k)}\xi_{\sigma(k+1)}-\xi_{\sigma(k+1)})+A_{\sigma^{(k)}}(\alpha+\beta\xi_{\sigma(k)}\xi_{\sigma(k+1)}-\xi_{\sigma(k)})\right)e^{-\sum_{j=1}^{N}(\alpha\xi_{i}^{-1}+\beta\xi_{i}-1)t}$
@@ 4.3.4 Bounds on the duality function
$=\prod_{j=1}^{l(S)}c_{\mathrm{p}}(1,\xi_{s_{j}},j-1,h_{s_{j}+1}^{+}(\xi),v,1,q)P_{\xi_{s_{j}}}(q^{-2},-vq^{h_{s_{j}+1}^{+}(\xi)+j-3},v^{-1}q^{h_{s_{j}+1}^{+}(\xi)-j+1},1;q^{2})$

### v2

@@ 2.3.3 Tracy–Widom contour integral formulas
$\begin{array}[]{l}\sum_{\sigma\in\mathbf{S}_{N}}\operatorname{sgn}\sigma\Big{(}\prod_{i<j}\left(\alpha+\beta\xi_{\sigma(i)}\xi_{\sigma(j)}-\xi_{\sigma(i)}\right)\\ \times\frac{\xi_{\sigma(2)}\xi_{\sigma(3)}^{2}\xi_{\sigma(4)}^{3}\cdots\xi_{\sigma(N)}^{N-1}}{\left(1-\xi_{\sigma(1)}\xi_{\sigma(2)}\xi_{\sigma(3)}\cdots\xi_{\sigma(N)}\right)\cdots\left(1-\xi_{\sigma(N-1)}\xi_{\sigma(N)}\right)\left(1-\xi_{\sigma(N)}\right)}\Big{)}=\alpha^{N(N-1)/2}\frac{\prod_{i<j}\left(\xi_{j}-\xi_{i}\right)}{\prod_{j}\left(1-\xi_{j}\right)}\end{array}$ (2)
@@ 2.5 From duality to asymptotics
$\lim_{L\rightarrow\infty}\sum_{S}\left(\sum_{x\in\mathcal{S}}\left|(n_{\lambda}^{S})^{-1/2}D_{\lambda}^{S}[x]-(n_{\overline{\lambda}}^{S})^{-1/2}D_{\overline{\lambda}}^{S}[x]\right|\cdot\left|Q_{\overline{\lambda}}(L,x)\right|\right)\\ \quad\quad\times\left(\sum_{y\in\mathcal{S}}\left|h(L)[y](n_{\lambda}^{S})^{-1/2}D_{\lambda}^{S}(y)W_{\lambda}(y)\right|\right)=0$ (8)
@@ 3.3 Particle locations in (multi–species) ASEP
×$\times\int_{C_{r}}\cdots\int_{C_{r}}A_{\sigma}\prod_{i=1}^{K}\xi_{\sigma(i)}^{x_{i}-y_{\sigma(i)}-1}\prod_{i=K+1}^{N}\frac{\xi_{\sigma(i)}^{M_{i}-y_{\sigma(i)}-1}}{1-\xi_{\sigma(i)}}e^{\sum_{i=1}^{N}(p\xi_{i}^{-1}+q\xi_{i}-1)t}d\xi_{1}\cdots d\xi_{N}.$ (11)
@@ 4.2.2 Proof of Proposition 2
$\sum_{\sigma\in S^{+}_{k}(N)}\left[\prod_{j=1}^{k-1}\xi_{\sigma(j)}^{\bar{x}_{j}-y_{\sigma(j)}-1}\right]\frac{\xi_{\sigma(k)}^{-M-y_{\sigma(k)}-1}\xi_{\sigma(k+1)}^{-M-y_{\sigma(k+1)}-1}}{(1-\xi_{\sigma(k)})(1-\xi_{\sigma(k+1)})}\left[\prod_{j=k+2}^{N}\xi_{\sigma(j)}^{\bar{x}_{j}-y_{\sigma(j)}-1}\right]\\ \times\alpha^{-1}\left(A_{\sigma}(\alpha+\beta\xi_{\sigma(k)}\xi_{\sigma(k+1)}-\xi_{\sigma(k+1)})+A_{\sigma^{(k)}}(\alpha+\beta\xi_{\sigma(k)}\xi_{\sigma(k+1)}-\xi_{\sigma(k)})\right)\\ \times e^{-\sum_{j=1}^{N}(\alpha\xi_{i}^{-1}+\beta\xi_{i}-1)t}$
@@ 4.3.4 Bounds on the duality function
$=\prod_{j=1}^{l(S)}c_{\mathrm{p}}(1,\xi_{s_{j}},j-1,h_{s_{j}+1}^{+}(\xi),v,1,q)$
×$\quad\quad\quad\times P_{\xi_{s_{j}}}(q^{-2},-vq^{h_{s_{j}+1}^{+}(\xi)+j-3},v^{-1}q^{h_{s_{j}+1}^{+}(\xi)-j+1},1;q^{2})$